# Topology-Driven Parallel Trajectory Optimization in Dynamic Environments

### v1

@@ Abstract
Ground robots navigating in complex, dynamic environments must compute collision-free trajectories to avoid obstacles safely and efficiently. Nonconvex optimization is a popular method to compute a trajectory in real-time. However, these methods often converge to locally optimal solutions and frequently switch between different local minima, leading to inefficient and unsafe robot motion. In this work, We propose a novel topology-driven trajectory optimization strategy for dynamic environments that plans multiple distinct evasive trajectories to enhance the robot’s behavior and efficiency. A global planner iteratively generates trajectories in distinct homotopy classes. These trajectories are then optimized by local planners working in parallel. While each planner shares the same navigation objectives, they are locally constrained to a specific homotopy class, meaning each local planner attempts a different evasive maneuver. The robot then executes the feasible trajectory with the lowest cost in a receding horizon manner. We demonstrate, on a mobile robot navigating among pedestrians, that our approach leads to faster and safer trajectories than existing planners.
@@ I INTRODUCTION
In this work, we present a global planner that identifies several distinct, high-level navigation options by considering the topology of the dynamic collision-free space. The underlying topology allows us to distinguish between similar and distinct trajectories. We then use each high-level trajectory as initialization for an optimization-based planner. The low-level planning problems are independent and are solved in parallel. Our framework does not modify the cost function of the optimization-based planner and can select the executed trajectory by comparing their optimal costs.
@@ II-2 Global Planners
In contrast with local planners, global planners do not rely on nonconvex optimization and therefore do not get trapped in local optima. Sampling-based global planners such as RRT* [7] and Probabilistic Roadmaps (PRM) [19] plan by randomly sampling and connecting states in the configuration space until a goal configuration is reached. Unfortunately, these methods become computationally intractable when dynamic constraints and collision avoidance are imposed on the problem and may return non-smooth trajectories.
Motion primitive planners (e.g., [6, 20]) instead generate a large number of trajectories that are dynamically feasible by construction. The best trajectory is identified by scoring each trajectory. Motion primitives planners efficiently compute smooth trajectories, but discretize the possible maneuvers which can lead to infeasibility and inefficient robot motion.
In [21], the motion planning problem with collision avoidance is solved via a convex optimization by utilizing Graphs of Convex Sets (GCS). This approach is promising, but is not real-time yet and imposes limitations on the trajectory end point, supported dynamics and constraints.
@@ II-3 Guidance Planners
To compute a suitable initial guess for a local planner considering obstacles, several authors [25, 26, 8, 11, 27, 28, 29] have noted that local optima related to collision avoidance link to the topology of trajectories through the collision-free space. Roughly speaking, two trajectories are in the same homotopy class if they can be smoothly transformed into each other in the collision-free space [26] (e.g., when they evade the obstacles on the same side).
Unfortunately, homotopy classes of trajectories are difficult to compute in general. If the environment can be consistently represented as a graph, then homotopy classes of trajectories can be identified through distinct paths over the graph [25]. This applies, for example, in structured autonomous driving applications through the lane structure of the road network [11, 27].
@@
Without structure in the environment, it is difficult to compute a trajectory in each homotopy class. Instead, several works, such as [29, 26, 8], compute distinct trajectories by filtering out homotopy equivalent trajectories during planning. For $3$-D navigation among static obstacles, [29] introduces Universal Visibility Deformation (UVD) to compare trajectories. Trajectories are UVD equivalent if they can be connected without collision at several intermediate times. The authors present a visibility-PRM [30] to generate UVD-distinct trajectories. Homotopy classes can be compared using homology classes as an approximation, as proposed by [26]. The work in [8], which relates most closely to this work, applied this approximation for $2$-D navigation among static obstacles. Their planner, Time Elastic Band (TEB) Local Planner, identifies several trajectories in distinct homology classes (using regular PRM) and uses each to initialize a soft-constrained optimization-based planner. TEB has, however, three main limitations that can hinder its performance in dynamic environments. Firstly, the trajectory topology is confined to the static workspace, treating dynamic obstacles and their future motion as static obstacles. Secondly, the guidance planner is designed to reach a single goal. Lastly, the local planner lacks hard constraints.
@@
In this work, we introduce a topology-guided planner that is different from these existing works in four ways. First, we consider homotopy classes in the dynamic collision-free space, that includes time, to incorporate the motion of dynamic obstacles (contrary to [8, 29]). Second, our framework does not modify the cost function (i.e., the performance criteria) of the local planner (contrary to [18, 21]). Third, our planner does not rely on a structured environment (contrary to [27, 11, 24]). Finally, our guidance planner can handle the case where its goal is blocked (contrary to [29, 8]) by considering multiple goal positions.
@@
In addition, we enforce the final trajectories to be in distinct homotopy classes using constraints in the local planner and we show that it is not sufficient to initialize the solver in a homotopy class (contrary to [8]). By consistently planning distinct trajectories, we can reidentify trajectories of prior planning iterations and use this information to make the planner more consistent and decisive.
@@ II-A Contribution
We validate our proposed framework in simulation on a mobile robot navigating among interactive pedestrians. We show how our proposed framework can accommodate different trajectory optimization approaches (e.g., [3] in the nominal case, and [4] to accommodate Gaussian uncertainties added to the motion of the dynamic obstacles). We show how our proposed framework enhances the performance of [3, 4] out of the box and we compare against three additional baselines ([6, 8, 28]). We finally demonstrate our planner in the real world on a mobile robot navigating among five pedestrians. Our C++/ROS implementation of T-MPC will be released open source.
@@ III PROBLEM FORMULATION
For dynamic collision avoidance, we consider the state space $\mathcal{X}\coloneqq\mathbb{R}^{2}\times[0,T]$, with $[0,T]$ a continuous finite time domain (see Fig. 1(b)). The area of the workspace occupied by the union of obstacles at time $t$ is denoted by $\mathcal{O}_{t}\subset\mathbb{R}^{2}$ and the obstacle set in the state space is thus $\mathcal{O}\coloneqq\bigcup_{\forall t\in[0,T]}(\mathcal{O}_{t},t)\subset\mathcal{X}$. The collision free state space (or free space) is denoted $\𝒪\mathcal{C}\coloneqq\mathcal{X}\textrm{\textbackslash}\mathcal{O}$. A trajectory is a continuous path through the state space, $\bm{\tau}:[0,1]\to\mathcal{X}$. The goal of the robot is to traverse along a given reference path $\bm{\gamma}:[0,1]\to\mathbb{R}^{2}$ without colliding with the obstacles. It is allowed to deviate from the path.
@@ III-A Optimization Problem
We formalize the planning problem as the following trajectory optimization problem over a horizon of $N_{L}$ steps
$\min_{\bm{u}\in\mathbb{U},\bm{x}\in\mathbb{X}}\quad$ $\sum_{k=0}^{N_{L}}J(\bm{x}_{k},\bm{u}_{k})$ (2a)
s.t. $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k}),\>\forall k,$ (2b)
@@ III-B Homotopic Trajectories
[31] (Homotopic Trajectories) Two paths connecting the same start and end points $\bm{x}_{s}$ and $\bm{x}_{g}$ respectively, are homotopic if they can be continuously deformed into each other without intersecting any obstacle. Formally, if $\bm{\tau}_{1},\bm{\tau}_{2}\in T$ represent two trajectories, with $\bm{\tau}_{1}(0)=\bm{\tau}_{2}(0)=\bm{x}_{s}$ and $\bm{\tau}_{1}(1)=\bm{\tau}_{2}(1)=\bm{x}_{g}$, then $\bm{\tau}_{1}$ is homotopic to $\bm{\tau}_{2}$ iff there exists a continuous map $\eta:[0,1]\times[0,1]\to\mathcal{C}$ such that $\eta(\alpha,0)=\tau_{1}(\alpha)\forall\alpha\in[0,1],\ \eta(\beta,1)=\tau_{2}(\beta),\forall\beta\in[0,1]$ and $\eta(0,\gamma)=\bm{x}_{s},\ \eta(1,\gamma)=\bm{x}_{s}\forall\gamma\in[0,1]$.
@@
$\mathcal{H}(\bm{\tau}_{i},\bm{\tau}_{j},\mathcal{O})=\begin{cases}1,&\bm{\tau}_{i},\bm{\tau}_{j}\textrm{ in the same homotopy class},\\ 0,&\textrm{otherwise}.\end{cases}$ (3)
Comparing homotopy classes is computationally inefficient. We therefore approximate homotopy classes through homology classes [31]. More details are provided in IV-B2. As the two definitions are almost always equivalent for mobile robot planning applications [8, 31], we will use the terminology homotopy throughout the rest of this paper.
@@ IV TOPOLOGY-DRIVEN MODEL PREDICTIVE CONTROL
where $\bm{x}_{0}$ denotes the robot initial state and $\mathcal{P}_{g}$ denotes a set of goal positions. Each local planner is initialized with one of the guidance trajectories and optimizes the trajectory in the same homotopy class. With $N_{G}$ and $N_{L}$ the horizon of the global and local planners, respectively, each local planner defines a mapping $L:\mathcal{X}^{N_{G}}\to\mathcal{X}^{N_{L}},$ $N_{L}\leq N_{G}$,
@@ IV-A Guidance Planner - Overview
The goal of the guidance planner is to quickly compute several homotopy distinct trajectories through the free space. Similarly to [29, 8], we perform this search via Visibility-Probabilistic RoadMaps (Visibility-PRM [30]), a sampling-based global planner. The modifications that we make ensure that the constructed graph remains consistent over successive iterations.
Figure 4: Illustration of the guidance planner in the state-space (time in the upwards axis). Visualization follows Fig. 3. (a) The visibility-PRM graph (black lines and dots) explores the free space toward the goals (orange dots). The homotopy distinct guidance paths (colored lines) are obtained by searching the graph. (b) The final trajectories are smoothened.
The guidance planner is outlined in Algorithm 1 and visualized in Fig. 4. Details of the algorithm are given in Sec. IV-B. We give a high-level overview here. First, Visibility-PRM constructs a sparse graph through the state space from the robot position to a set of goals, where each connection is homotopy distinct (line $1$). The goals represent end points for the guidance planner and are placed along the reference path. For each goal, DepthFirstSearch searches in this graph for the shortest $P$ paths that reach it (line $2$). Any homotopy equivalent trajectories are filtered out by FilterAndSelect (line $3$), ensuring that the remaining trajectories are in distinct homotopy classes. The $P$ trajectories that seem most promising are filtered out based on a heuristic that prefers its end point to be as close as possible to the reference path at the reference velocity (see Fig. 3(a)). In IdentifyAndPropagate (line $4$), the resulting $P$ trajectories are compared to previously found trajectories to track their homotopy class over consecutive iterations of the guidance planner. Finally, nodes in the Visibility-PRM graph are propagated by dropping their time index by the planning time.
Through this process, we obtain in each iteration $P$ piecewise linear trajectories $\mathcal{T}_{P}=\{\bm{\tau}_{1},\ldots,\bm{\tau}_{P}\}$ that each connects the robot position to one of the goals (see Fig. 3). We finally smoothen these trajectories and fit cubic splines to make them differentiable (see Fig. 3(b)). More details can be found in [28].
@@
Input: $\mathcal{C}$, $\bm{x}_{0}$, $\bm{x}_{N}$, previous nodes $\mathcal{G}^{-}$
1 $\mathcal{G}\ \leftarrow$ Visibility-PRM$(\mathcal{C},\bm{x}_{0},\bm{x}_{N},\mathcal{G}^{-})$
@@ IV-B Guidance Planner - Detailed Description
In the following, we discuss our Visibility-PRM implementation of the guidance planner and how we distinguish between trajectories in distinct homotopy classes.
IV-B1 Multiple Goals in Visibility-PRM
@@
The visibility-PRM scheme that we use is a direct extension of [28], we refer to [28, Algorithm 1] for the basic algorithm. In this work, we address the limitation that a single goal must be reached by the planner that causes the planner to fail frequently. Our modification supports multiple goals to make the planner more robust. We propose a new “Goal” node type to Visibility-PRM next to the existing Guard (cannot connect to any other nodes) and Connector (connects to exactly two Guards) nodes. Goals inherit the properties of Guards but are inserted initially and are likely visible to each other. When a Connector can connect to multiple Goals, we single out the Goal with the lowest distance to the point on the reference path reached with the reference velocity (i.e., our ideal goal). This generalization of the goal representation makes the guidance planner more robust. In our implementation, we deploy a grid of goals centered around the reference path (see Fig. 4).
IV-B2 H-signature
We use the H-signature [31] to compare homology classes of trajectories, approximating their homotopy classes.
Definition 2.
If two trajectories are homotopic, they are homologous. The reverse does not hold.
Figure 5: Skeletons used for computing the H-signature [31]. Blue link $2$ follows the motion prediction of the pedestrian. Blue link $3$ moves outside of the environment at some $t>T$ linking to $1$ that is placed at $t<0$. The loops formed by the homology distinct trajectories each enclose the skeletons in a unique way.
To compute the H-signature, the obstacles and trajectories need to be processed. For each obstacle, a closed-loop must be constructed that describes its motion over time, known as a skeleton (see Fig. 5). We construct the skeletons by following the motion predictions of the obstacles for $t=0$ to $t=T$. Then, on either side of the time axis, the trajectories are extended out of the planning horizon and finally, the skeleton is closed outside of the practical environment. The two trajectories to be compared are merged into a loop. Because we consider multiple goals, the end points of the trajectories may not be the same. In this case, we connect them directly at $t=T$.
We assume that obstacle trajectories are piecewise linear (e.g., discrete-time trajectories). Denoting linear segment $j$ of the obstacle skeleton as $\overline{s_{j}s_{j}^{\prime}}$, the H-signature contribution for position $\bm{r}$ along the loop of the trajectories for each obstacle is computed by (see [31])
$\bm{B}(\bm{r})$ $=\frac{1}{4\pi}\sum_{j=1}^{n_{i}}\bm{\Phi}(\bm{s}_{j},\bm{s}_{j}^{\prime},\bm{r})d\bm{r},$ (6)
$\bm{\Phi}(\bm{s}_{j},\bm{s}_{j}^{\prime},\bm{r})$ $=\frac{1}{||\bm{d}||^{2}}\left(\frac{\bm{d}\times\bm{p}^{\prime}}{||\bm{p}^{\prime}||}-\frac{\bm{d}\times\bm{p}}{||\bm{p}||}\right),$ (7)
$\bm{p}=\bm{s}_{j}-\bm{r}$ $,\ \bm{p}^{\prime}=\bm{s}_{j}^{\prime}-\bm{r},\ \bm{d}=\frac{(\bm{s}_{j}^{\prime}-\bm{s}_{j})\times(\bm{p}\times\bm{p}^{\prime})}{||\bm{s}_{j}^{\prime}-\bm{s}_{j}||^{2}}$ (8)
Integrating over the looped trajectories $\bm{l}$ gives
[…]
which equals $1$ if the obstacles is enclosed and 00 otherwise. We use the GSL library [32] to perform the integration and cache the H-signature for each trajectory to prevent re-computation. Topology comparison function (3) is implemented by computing the integral for all obstacles, returning Distinct if one of the integrals equals $1$ and Equivalent otherwise.

### v2

@@ Abstract
Ground robots navigating in complex, dynamic environments must compute collision-free trajectories to avoid obstacles safely and efficiently. Nonconvex optimization is a popular method to compute a trajectory in real-time. However, these methods often converge to locally optimal solutions and frequently switch between different local minima, leading to inefficient and unsafe robot motion. In this work, we propose a novel topology-driven trajectory optimization strategy for dynamic environments that plans multiple distinct evasive trajectories to enhance the robot’s behavior and efficiency. A global planner iteratively generates trajectories in distinct homotopy classes. These trajectories are then optimized by local planners working in parallel. While each planner shares the same navigation objectives, they are locally constrained to a specific homotopy class, meaning each local planner attempts a different evasive maneuver. The robot then executes the feasible trajectory with the lowest cost in a receding horizon manner. We demonstrate, on a mobile robot navigating among pedestrians, that our approach leads to faster trajectories than existing planners.
@@ I INTRODUCTION
We present a global planner that identifies several distinct, high-level navigation options by considering the topology of the dynamic collision-free space. The underlying topology allows us to distinguish between similar and distinct trajectories. We then use each high-level trajectory as initialization for an optimization-based planner. The low-level planning problems are independent and are solved in parallel. Our framework does not modify the cost function of the optimization-based planner and can select the executed trajectory by comparing their optimal costs.
@@ II-2 Global Planners
[…]
Motion primitive planners (e.g., [6, 22]) instead generate a large number of trajectories that are dynamically feasible by construction. The best trajectory is identified by scoring each trajectory. Motion primitives planners efficiently compute smooth trajectories, but discretize the possible maneuvers which can lead to infeasibility and inefficient robot motion. For static environments, [23] presents a smooth global planner that repairs dynamic mismatches between global plans through trajectory optimization. It is however computationally demanding. Similarly, PiP-X [24] combines graph-search with funnels to find robust dynamically feasible paths but may return inefficient trajectories. In [25], the motion planning problem with collision avoidance is solved via a convex optimization by utilizing Graphs of Convex Sets (GCS). This approach is promising, but is not real-time yet and imposes limitations on the trajectory end point, supported dynamics and constraints.
@@ II-3 Guidance Planners
To compute a suitable initial guess for a local planner considering obstacles, several authors [29, 30, 8, 11, 31, 32, 33, 34, 35] have noted that local optima related to collision avoidance link to the topology of trajectories through the collision-free space. Roughly speaking, two trajectories are in the same homotopy class if they can be smoothly transformed into each other in the collision-free space [30] (e.g., when they evade the obstacles on the same side). Unfortunately, homotopy classes of trajectories are difficult to compute in general. If the environment can be consistently represented as a graph, then homotopy classes of trajectories can be identified through distinct paths over the graph [29]. This applies, for example, in structured autonomous driving applications through the lane structure of the road network [11, 31].
Without structure in the environment, it is difficult to compute a trajectory in each homotopy class. Graphs can be constructed from static obstacles. In [36], Delauney triangulation is used to identify passable gaps between dynamic obstacles in a global planner. Voronoi graphs are used in [34] to identify homotopy classes with respect to static obstacles and in [8] include each dynamic obstacle and their predicted motion as a static obstacle. Trajectories are generated from the homotopy class description in [37] by modeling interactions as a physical vortex system. These graph-based and generative approaches can exhaust the possible homotopy classes, but scale poorly to crowded environments (as noted in [8] and [37]).
@@
Instead, several works, such as [33, 30, 8], compute distinct trajectories by filtering out homotopy equivalent trajectories during planning. For $3$-D navigation among static obstacles, [33] introduces Universal Visibility Deformation (UVD) to compare trajectories. Trajectories are UVD equivalent if they can be connected without collision at several intermediate times. The authors present a visibility-PRM [38] to generate UVD-distinct trajectories. In $2$D dynamic environments, homotopy classes are typically compared via winding numbers [39] or the H-signature [30]. Winding numbers track the relative angle between the robot and dynamic obstacles over their trajectories. They were used in [34] to distinguish homotopy classes of trajectories with respect to dynamic obstacles. In [35], an MPC with winding numbers in the cost function was proposed to motivate passing progress. The H-signature uses homology classes as an approximation for homotopy classes. The work in [8], which relates most closely to this work, applied this approximation for $2$-D navigation among static obstacles. Their planner, Time Elastic Band (TEB) Local Planner, identifies several trajectories in distinct homology classes (using regular PRM) and uses each to initialize a soft-constrained optimization-based planner. TEB has, however, three main limitations that can hinder its performance in dynamic environments. Firstly, the trajectory topology is confined to the static workspace, treating dynamic obstacles and their future motion as static obstacles. Secondly, the guidance planner is designed to reach a single goal. Lastly, the local planner lacks hard constraints.
@@
In this work, we introduce a topology-guided planner that is different from these existing works in four ways. First, we consider homotopy classes in the dynamic collision-free space, that includes time, to incorporate the motion of dynamic obstacles (contrary to [8, 33]). Second, our framework does not modify the cost function (i.e., the performance criteria) of the local planner (contrary to [35, 18, 25]). Third, our planner does not rely on a structured environment (contrary to [31, 11, 28]). Finally, our guidance planner can handle the case where its goal is blocked (contrary to [33, 8]) by considering multiple goal positions.
@@
In addition, we enforce the final trajectories to be in distinct homotopy classes using constraints in the local planner and we show that it is not sufficient to initialize the solver in a homotopy class (contrary to [8]). By consistently planning distinct trajectories, we can reidentify trajectories of prior planning iterations and use this information to make the planner more consistent and decisive. Our method furthermore supports the H-signature, winding numbers and UVD for comparing homotopy classes.
@@ II-A Contribution
We validate our proposed framework in simulation on a mobile robot navigating among interactive pedestrians. We show how our framework can accommodate different trajectory optimization approaches (e.g., [3] in the nominal case, and [4] to accommodate Gaussian uncertainties added to the motion of the dynamic obstacles). We show how our framework enhances the performance of [3, 4] out of the box and we compare against three additional baselines ([6, 8, 32]). We finally demonstrate our planner in the real world on a mobile robot navigating among five pedestrians. Our C++/ROS implementation of T-MPC will be released open source.
@@ III PROBLEM FORMULATION
For high-level planning with dynamic collision avoidance, we consider the simplified state space $\mathcal{X}\coloneqq\mathbb{R}^{2}\times[0,T]$, with $[0,T]$ a continuous finite time domain (see Fig. 2(b)). The area of the workspace occupied by the union of obstacles at time $t$ is denoted by $\mathcal{O}_{t}\subset\mathbb{R}^{2}$ and the obstacle set in the state space is thus $\mathcal{O}\coloneqq\bigcup_{\forall t\in[0,T]}(\mathcal{O}_{t},t)\subset\mathcal{X}$. The collision free state space (or free space) is denoted $\𝒪\mathcal{C}\coloneqq\mathcal{X}\textrm{\textbackslash}\mathcal{O}$. A trajectory is a continuous path through the state space, $\bm{\tau}:[0,1]\to\mathcal{X}$. The goal of the robot is to traverse along a given reference path $\bm{\gamma}:[0,1]\to\mathbb{R}^{2}$ without colliding with the obstacles while tracking a reference velocity $v_{\textrm{ref}}$. It is allowed to deviate from the path.
@@ III-A Optimization Problem
We formalize the planning problem as the following trajectory optimization problem over a horizon of $N$ steps
$\min_{\bm{u}\in\mathbb{U},\bm{x}\in\mathbb{X}}\quad$ $\sum_{k=0}^{N}J(\bm{x}_{k},\bm{u}_{k})$ (2a)
s.t. $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k}),\>\forall k$ (2b)
@@ III-B Homotopic Trajectories
[30] (Homotopic Trajectories) Two paths connecting the same start and end points $\bm{x}_{s}$ and $\bm{x}_{g}$ respectively, are homotopic if they can be continuously deformed into each other without intersecting any obstacle. Formally, if $\bm{\tau}_{1},\bm{\tau}_{2}\in T$ represent two trajectories, with $\bm{\tau}_{1}(0)=\bm{\tau}_{2}(0)=\bm{x}_{s}$ and $\bm{\tau}_{1}(1)=\bm{\tau}_{2}(1)=\bm{x}_{g}$, then $\bm{\tau}_{1}$ is homotopic to $\bm{\tau}_{2}$ iff there exists a continuous map $\eta:[0,1]\times[0,1]\to\mathcal{C}$ such that $\eta(\alpha,0)=\tau_{1}(\alpha)\forall\alpha\in[0,1],\ \eta(\beta,1)=\tau_{2}(\beta),\forall\beta\in[0,1]$ and $\eta(0,\gamma)=\bm{x}_{s},\ \eta(1,\gamma)=\bm{x}_{g}\forall\gamma\in[0,1]$.
@@
$\mathcal{H}(\bm{\tau}_{i},\bm{\tau}_{j},\mathcal{O})\!=\!\begin{cases}1,&\bm{\tau}_{i},\bm{\tau}_{j}\textrm{ in the same homotopy class}\\ 0,&\textrm{otherwise}\end{cases}$ (3)
Verifying whether two trajectories are in the same homotopy class can be computationally inefficient. We support the H-signature [30], winding numbers [39] and UVD [33] that allow us to approximately perform this verification in real-time. Details of the three methods are provided in Appendix A.
@@ IV TOPOLOGY-DRIVEN MODEL PREDICTIVE CONTROL
where $\bm{x}_{0}$ denotes the robot initial state and $\mathcal{P}_{g}$ denotes a set of goal positions. Each local planner is initialized with one of the guidance trajectories and optimizes the trajectory in the same homotopy class. With $N$ the horizon of the guidance and local planners [^1] The guidance planner horizon could extend beyond the horizon of the local planner. We set them equal for simplicity., each local planner defines a mapping $L:\mathcal{X}^{N}\to\mathcal{X}^{N}$,
@@ IV-A Guidance Planner - Overview
The goal of the guidance planner is to quickly compute several homotopy distinct trajectories through the free space. Similarly to [33, 8], we perform this search via Visibility-Probabilistic RoadMaps (Visibility-PRM [38]), a sampling-based global planner. The modifications that we make ensure that the graph remains consistent over successive iterations.
Figure 4: Illustration of the guidance planner in the state-space (time in the upwards axis). Visualization follows Fig. 3. (a) The visibility-PRM graph (black lines and dots) explores the free space toward the goals placed at $t=T$ around the reference path (orange dots). The homotopy distinct guidance paths (colored lines) are obtained by searching the graph. (b) The final trajectories are smoothened.
The guidance planner is outlined in Algorithm 1 and visualized in Fig. 4. Details of the algorithm are given in Sec. IV-B. We give a high-level overview here. First, Visibility-PRM constructs a sparse graph through the state space from the robot position to a set of goals, where each connection is homotopy distinct (line $1$). The goals represent end points for the guidance planner and are placed along the reference path. For each goal, DepthFirstSearch (line $2$) searches in this graph for the shortest $P$ trajectories that reach it. Any homotopy equivalent trajectories are filtered out by FilterAndSelect (line $3$), ensuring that the remaining trajectories are in distinct homotopy classes. The $P$ trajectories that seem most promising are selected by a heuristic that prefers its goal to be as close as possible to the reference path at the reference velocity (see Fig. 4(a)). IdentifyAndPropagate (line $4$) verifies if any of the selected trajectories are equivalent to trajectories of the previous planning iteration. This reidentification makes it possible to follow the same passing behavior over multiple planning iterations. We finally propagate the nodes in the Visibility-PRM graph by lowering their time state by the planning time step.
Through this process, we obtain in each iteration $P$ piecewise linear trajectories $\mathcal{T}_{P}=\{\bm{\tau}_{1},\ldots,\bm{\tau}_{P}\}$ that each connects the robot position to one of the goals (see Fig. 3). We finally smoothen these trajectories and fit cubic splines to make them differentiable (see Fig. 4(b)). More details can be found in [32]. The smoothening procedure produces only a small displacement in trajectories to maintain their homotopy class.
@@
Input: $\mathcal{C}$, $\bm{x}_{0}$, $\bm{x}_{N}$, previous graph $\mathcal{G}^{-}$, previous trajectories $\mathcal{T}_{P}^{-}$
1
[…]
@@ IV-B Guidance Planner - Detailed Description
We detail each step of Algorithm 1 in the following.
Visibility-PRM computes sparse paths through the free space by randomly sampling positions and creating either a Guard or Connector node at the sampled position. The type of node depends on the number of Guards that it can directly connect to without colliding (i.e., which Guards are visible). A Guard is added if no other Guards are visible. A Connector (see black dots in Fig. 4) is added when exactly two Guards are visible and its connection to the Guards is feasible (e.g., satisfying velocity and acceleration limits). Similar to [33], we also check if any Connectors link to the same Guards (referred to as neighbors). If there are neighbors, we keep the new connection if it is distinct from existing connections, which we verify with homotopy comparison function (3). If it is equivalent and more efficient than the existing connection (e.g., if its connection is shorter), then we replace the existing connector with the new connector. In regular visibility-PRM [38], the graph is initialized with a Guard at the start and goal positions and new nodes are drawn up to a time or node limit. More details of the algorithm can be found in [32, Algorithm 1].
Multiple Goals in Visibility-PRM
@@
In this work, we address the limitation that a single goal must be reached by Visibility-PRM, which causes the planner to fail when that goal cannot be reached. We propose to add a Goal node type to Visibility-PRM. Goals inherit the properties of Guards but are inserted initially and are likely visible to each other. When a Connector can connect to multiple Goals, we single out the Goal with the lowest distance to the point on the reference path reached with the reference velocity (i.e., our ideal goal). By supporting multiple goals, we increase the robustness of the guidance planner. In practice, we deploy a grid of goals centered around the reference path (see Fig. 4).
Homotopy Comparison
We use the homotopy comparison function (3) to verify if two trajectories are in the same homotopy class. We implemented (3) with the H-signature [30], winding numbers [39] and UVD [33]. Appendix A provides details on these methods. In our experiments, we use the H-signature that joins the two trajectories to be compared into a loop and verifies if that loop encircles any moving obstacles. If it does, then the two trajectories pass obstacles differently and belong to different homotopy classes.
DepthFirstSearch searches for $P$ paths to each goal, with each search implemented similar to [8, Algorithm 1].
FilterAndSelect uses homotopy comparison function (3) to remove equivalent trajectories to different Goals found by DepthFirstSearch. The set of filtered trajectories $\mathcal{T}_{F}$ therefore satisfy
$\mathcal{H}(\bm{\tau}_{i},\bm{\tau}_{j},\mathcal{O})=0,\ \forall i,j,i\neq j,\ \bm{\tau}_{i},\bm{\tau}_{j}\in\mathcal{T}_{F}.$ (6)
The $P$ lowest cost trajectories in $\mathcal{T}_{F}$ constitute the output $\mathcal{T}_{P}$.
IdentifyAndPropagate uses homotopy comparison function (3) to link new trajectories to trajectories found in the previous iteration. It checks for each previous trajectory $\bm{\tau}_{i}^{-}\in\mathcal{T}^{-}_{P}$ if
$\exists\bm{\tau}_{j}\in\mathcal{T}_{P},\mathcal{H}(\bm{\tau}_{i}^{-},\bm{\tau}_{j})=1.$ (7)
A unique identifier, tied to the homotopy class, is passed from $\bm{\tau}_{i}^{-}$ to $\bm{\tau}_{j}$ if the latter exists. We can use this identifier to decide which trajectory to follow (see Sec. IV-E).